FTIR SPECTROSCOPY OF NO_{3} : OBSERVATION AND ANALYSIS OF THE 1127 cm^{−1} BAND.

\underline{K. KAWAGUCHI}, N. SHIMIZU, R. FUJIMORI, Department of Chemistry, Faculty of Science, Okayama University, 3-1-1, Tsushima-Naka, Okayama 700-8530, JAPAN; T. ISHIWATA, Graduate School of Information Sciences, Hiroshima City University, 3-4-1 Otsuka-Higashi, Hiroshima 731-3194, JAPAN; I. TANAKA, Tokyo Institute of Technology, 2-12-1, Ookayama, Meguro, Tokyo 152-8550, Japan; and E. HIROTA, The Graduate University for Advanced Studies, Hayama, Kanagawa 240-0193, JAPAN.

We report the infrared spectrum of NO_{3} studied with a Bruker 120 HR. In the present experiment, the NO_{3} production was monitored by using He-Ne laser absorption at 633 nm. The multi-path cell was also improved to attain 48 m effective path length, and a new HgCdTe detector was introduced. In the 1000-1900 cm^{−1} region, the following 7 bands were observed at 1127, 1137, 1173, 1413, 1492, 1550, 1927 cm^{−1}. The 1127, 1137 and 1550 bands were not reported by matrix isolation, and these are thought to be due to hot bands from the ν_{4} state. The analysis of the 1127 band was carried out by fixing the upper state parameters to those of the 1492 band, and major molecular constants are determined as follows, B=0.459339(12), C=0.228848(20), Cζ_{4}=-0.044279(29), ν_{4}=365.4887(13) cm^{−1}. The two A-E type bands are also assigned in the 1133 and 1137 cm^{−1} regions in addition to E-E type, indicating that the 1492 band is assignable to ν_{3} + ν_{4}, in contrast to the previous ν_{3} assignment.